When it say find the x intercepts do I graph it or give the variable for the intercept
y=x^2-9x+20





Answer by funmath(2933)

When it say find the x intercepts do I graph it or give the variable for the intercept
y=x^2-9x+20
Usually, if they don't give you the graph or ask for the graph, they want the coordinate of the x intercepts. We find x-intercepts by letting y=0 and solveing for x.
0=x^2-9x+20 factor
o=(x__)(x__)
fille the blanks with two numbers that multiply to give you +20, but add to give you -9. -4*-5=+20 and -4+(-5)=-9
0=(x-4)(x-5) Let each parentheses=0 and solve for x.
x-4=0 and x-5=0
x-4+4=0+4 and x-5+5=0+5
x=4 and x=5
The x-intercept are (4,0) and (5,0)
Happy Calculating and Happy Thanksgiving!!!